# Exploring Exponential Runge-Kutta Methods: A Survey

### v1

@@ 5. Examples
To conclude our manuscript, we present two low-dimensional examples useful for demonstrating the functionalities of second-order ExpRK method (ExpRK2) and their comparison with other numerical methods of equal order, namely the standard second-order RK method (RK2) and the classic second-order Rosenbrock method (Rb2). In this way, we aim to provide both a more complete description of the numerical methods analyzed in this manuscript and a useful tool for approaching ExpRK methods. We remark that, in 2020, Montanelli and Bootland [76] published a comprehensive review of exponential methods, comparing various formulas for solving stiff PDEs in multiple dimensions. Their analysis, focused on periodic boundary conditions and semilinear stiff PDEs, concluded that the ETDRK4 scheme by Cox and Matthews [19] remains one of the most effective approaches despite the proliferation of more complex alternatives. Their work is particularly notable for its examination of both diffusive problems (characterized by large negative eigenvalues) and dispersive problems (distinguished by their highly oscillatory nature), as well as its discussion of computational techniques for the $\varphi$-functions based on contour integrals.
The two didactic examples we propose are one-dimensional and two-dimensional (where here the ‘dimension’ refers to the size of the system of ODEs), respectively. Specifically, for the 1D case, we revisit the introductory example presented by Cox and Matthews in 2002 [19, Sect. 4.1]; and for the 2D case, we examine a classical model from nonlinear mechanics and dynamical systems, namely the Duffing oscillator [^12] For a comprehensive analysis of the nonlinear Duffing oscillator and its applications, we recommend the following volume [53] to the interested reader.. In what follows, we briefly three second-order methods for solving stiff systems of ODEs. As already noticed, stiff ODEs require specialized numerical techniques due to their rapidly changing components existing alongside slowly evolving ones. The three methods under consideration–RK2, ExpRK2 and Rb2–represent different approaches to handling stiffness, while all maintaining second-order accuracy. Their properties suggest different performance characteristics when applied to stiff problems.
@@
In the first example we present, there exists an analytical expression for the solution $u$ of the one-dimensional ODE, and we use this exact expression in the formula above. In the second two-dimensional example, such an explicit, exact expression does not exist, and we compare the methods against a reference expression for the solution $u$. Additionally, we measure the computational time required by each method, enabling a comprehensive comparison of both accuracy and efficiency. This dual assessment, as emphasized earlier, is particularly important when dealing with stiff systems, where the choice of numerical method significantly impacts the simulation’s precision and performance. Through this analysis, we aim to further illustrate the practical considerations discussed in previous sections and provide additional insights into the behavior of these methods across different problem types and stiffness regimes.
The Three Methods
Here, we describe the three numerical methods which we have used for our simulations, namely for solving (in one case non-autonomous, in one case autonomous) systems of ODEs of the form:
@@
$\mathbf{a}$ $=e^{hA}\mathbf{u}_{n}+h\varphi_{1}(hA)\mathbf{g}(t_{n},\mathbf{u}_{n})\,$
@@
is the Jacobian matrix, and $\gamma$ is a stabilization parameter.
@@
Following the same approach presented by Cox and Matthews [19] in their analysis of ETD methods, we consider a non-autonomous ODE model with rapid decay, characteristic of stiff systems:
@@
where we have chosen $k=-100$ to generate the rapid linear decay typical of stiff problems, in contrast with the $\mathcal{O}(1)$ time scale of the forcing term. This equation, while simple in form, captures the essential features of stiff systems: a fast decaying homogeneous solution combined with a (relatively) slow varying forcing term. For subsequent evaluation of the different numerical schemes, we utilize the exact analytical solution given by:
@@
The behavior of this differential equation can be divided into two distinct phases: an initial fast transient phase, during which the solution rapidly approaches the slow manifold, followed by a phase in which the solution evolves along this slow manifold. As noticed in [19], when $k<0$ and $|k|\gg 1$, the solution quickly approaches a slow manifold on which
@@
A robust numerical method must accurately capture both of these phases and work effectively when the time step $h$ is of the order of $1/|k|$. This test case allows us to evaluate how well our three second-order methods handle the challenges posed by stiffness, particularly the balance between stability and accuracy when using relatively large time steps compared to the fastest time scale in the system. We examined the performance of the methods over the interval $[0,\pi/2]$ using various time steps ranging from $h=0.0001$ to $h=0.1$, comparing both the accuracy of the numerical solutions against the analytical solution and the computational efficiency of each method.
Table 2. Numerical methods performance for various step sizes
@@
$1\times 10^{-4}$ $3.5469\times 10^{-8}$ $3.5892\times 10^{-8}$ $4.6309\times 10^{-7}$ 0.20067 0.15224 0.13453
$5\times 10^{-4}$ $1.9839\times 10^{-6}$ $1.9952\times 10^{-6}$ $4.3310\times 10^{-7}$ 0.03203 0.03247 0.03199
$1\times 10^{-3}$ $1.8844\times 10^{-6}$ $1.9327\times 10^{-6}$ $2.8822\times 10^{-6}$ 0.01465 0.01640 0.00999
$5\times 10^{-3}$ $1.2446\times 10^{-5}$ $1.0355\times 10^{-5}$ $3.5272\times 10^{-5}$ 0.00367 0.00357 0.00250
$1\times 10^{-2}$ $3.3266\times 10^{-5}$ $1.6478\times 10^{-5}$ $6.2265\times 10^{-5}$ 0.00338 0.00402 0.00365
$5\times 10^{-2}$ $6.4874\times 10^{30}$ $5.7744\times 10^{-4}$ $1.1943\times 10^{-3}$ 0.00024 0.00018 0.00028
$1\times 10^{-1}$ $6.3772\times 10^{27}$ $5.3437\times 10^{-4}$ $1.5318\times 10^{-1}$ 0.00007 0.00014 0.00022
@@
Figure 1. Relative error as a function of computational time, highlighting the efficiency of each method (1(a)) and relative error as a function of step size, showing how accuracy evolves with different step choices (1(b)).
@@
The results presented in Figure 1 and Table 2 highlight the distinct behaviors of the three second-order numerical methods under analysis.
@@
RK2 demonstrates stability only for very small step sizes, specifically for values of $h\leq 0.001$. However, for $h\geq 0.05$, the relative error grows uncontrollably. This behavior confirms a strict stability limit around $h=0.001$, beyond which the method fails to provide reliable results. On the other hand, ExpRK2 exhibits excellent stability across all tested step sizes. Even for larger ones, the method maintains a reasonably small relative error. Rb2 also demonstrates stable behavior, albeit within a more restricted range. For step sizes up to $0.01$, the method produces accurate results. However, for $h=0.05$, the relative error starts increasing significantly, and for $h=0.1$ it becomes quite large ($1.53\times 10^{-1}$). The superior accuracy makes ExpRK2 a strong candidate for solving stiff problems.
Regarding computational efficiency, Rb2 is the fastest for very small time steps such as $h=0.0001$, completing computations in approximately 0.134 seconds. ExpRK2 follows closely at 0.152 seconds, while RK2 is the slowest at 0.201 seconds. For step sizes in the range $0.001\leq h\leq 0.01$, all three methods exhibit similar performance. For large step sizes ($h\geq 0.05$), computation times become negligible for all methods, remaining below 0.0002 seconds, making observations about performance insignificant.
@@
Considering both accuracy and efficiency, ExpRK2 emerges as the best-balanced method. It sustains high accuracy even with large step sizes, allowing for significantly larger steps compared to the other methods. Although it has a slight computational overhead per step, the reduced number of required steps makes it more efficient overall.
The Rosenbrock method offers slightly faster computation times for very small step sizes but with stricter stability constraints: the method appears to work very effectively for a problem of small scale. RK2, in contrast, proves to be the least suitable method for this stiff problem, as it requires extremely small time steps to maintain stability and lacks competitiveness in both accuracy and computational efficiency.
These findings align with the conclusions drawn by Cox and Matthews in [19], confirming the superiority of exponential methods for solving stiff problems. In particular, the ExpRK2 method offers a significant advantage in terms of stability and accuracy, permitting the use of much larger step sizes than conventional methods. For stiff problems such as the one studied here, characterized by a parameter $k=-100$, ExpRK2 represents the optimal choice among the three second-order methods, striking an ideal balance between accuracy, stability, and computational efficiency.
@@
Table 3. Numerical methods performance for various step sizes.
$h$ RK2 Error ExpRK Error Rb2 Error RK2 Time (s) ExpRK Time (s) Rb2 Time (s)
$1\times 10^{-4}$ $3.7937\times 10^{-5}$ $3.0147\times 10^{-5}$ $1.1169\times 10^{-4}$ 0.109362 21.442708 1.041663
$5\times 10^{-4}$ $1.9122\times 10^{-3}$ $2.8181\times 10^{-4}$ $2.3210\times 10^{-3}$ 0.017562 4.002774 0.143251
$1\times 10^{-3}$ $9.6389\times 10^{-3}$ $1.5939\times 10^{-3}$ $9.7571\times 10^{-3}$ 0.008851 2.099005 0.073638
$5\times 10^{-3}$ $6.3872\times 10^{-1}$ $1.4550\times 10^{-1}$ $3.4412\times 10^{-1}$ 0.002555 0.431901 0.019036
$1\times 10^{-2}$ $2.9671\times 10^{0}$ $9.5181\times 10^{-1}$ $1.1835\times 10^{0}$ 0.005627 0.238445 0.010617
$5\times 10^{-2}$ NaN $1.3049\times 10^{0}$ $1.2755\times 10^{-1}$ 0.000092 0.086358 0.003250
$1\times 10^{-1}$ NaN $1.6391\times 10^{0}$ $8.3403\times 10^{-1}$ 0.000030 0.033348 0.001458
Figure 2. Relative error as a function of step size, showing how accuracy evolves with different step choices (2(a)) and relative error as a function of computational time, highlighting the efficiency of each method (2(b)).
The results presented in Figure 2 and Table 3 provide a comprehensive analysis of three second-order numerical methods applied to the Duffing oscillator system [^13] All relative errors are computed with respect to a high-precision numerical solution obtained using MATLAB’s ode15s solver, as no analytical solution is available for this nonlinear system. with $\omega=1$ and $k=100$.
RK2 demonstrates severe stability limitations when applied to this stiff problem. For step sizes $h\leq 0.001$, the method produces reasonable errors, but as $h$ increases to $0.005$, the relative error grows to approximately $6.39\times 10^{-1}$. At $h=0.01$, the error reaches $2.97$, indicating complete loss of accuracy. For larger step sizes ($h\geq 0.05$), the method fails entirely, producing NaN values. This behavior confirms the strict stability constraints of explicit Runge-Kutta methods when applied to stiff problems.
@@
ExpRK2 demonstrates superior accuracy for most step sizes up to $h=0.01$, particularly excelling at moderate step sizes where it maintains errors significantly lower than the other methods. For $h=5\times 10^{-4}$ and $h=1\times 10^{-3}$, ExpRK2 shows errors approximately an order of magnitude smaller than both RK2 and Rb2, despite not being designed so as to capture stiffness phenomena caused by the nonlinear terms. However, for larger step sizes ($h\geq 0.05$), the errors increase substantially, exceeding the error of Rb2 at these larger step sizes.
Regarding computational efficiency, the picture is quite different from the previous analysis. RK2 is consistently the fastest method across all step sizes, with computation times decreasing from approximately $0.11$ seconds at $h=0.0001$ to just $0.00003$ seconds at $h=0.1$. The Rosenbrock method occupies an intermediate position, with computation times ranging from $1.04$ seconds to $0.00146$ seconds. Notably, ExpRK2 emerges as the most computationally intensive method by a substantial margin, requiring $21.44$ seconds for the smallest step size and $0.033$ seconds for the largest one. This represents a substantial computational overhead compared to the other methods, over 20 times slower than Rb2 for the smallest step size.
@@
The analysis of these numerical methods for the Duffing oscillator reveals important trade-offs between accuracy, stability, and computational efficiency. RK2, while computationally inexpensive, suffers from severe stability limitations that render it unsuitable for step sizes larger than $0.001$. The Rosenbrock method offers an attractive balance, providing reasonable stability across a wide range of step sizes while maintaining moderate computational costs. Its non-monotonic error behavior at larger step sizes warrants further investigation but could potentially be exploited in specific applications.
ExpRK2 demonstrates excellent accuracy, particularly at moderate step sizes where it significantly outperforms both RK2 and Rb2. However, it incurs substantially higher computational costs than both alternative methods. The extreme computational overhead of ExpRK2 in this implementation—more than 20 times slower than Rb2 at the smallest step size—suggests that the method may require significant optimization before practical application to time-sensitive problems.
For the Duffing oscillator with the given parameters, the Rb2 appears to offer the most favorable combination of stability, accuracy, and computational efficiency. It maintains acceptable accuracy even at large step sizes and does so with reasonable computational cost. This makes it particularly suitable for applications where a balance between accuracy and efficiency is required. However, if highest accuracy is the primary concern and computational resources are abundant, ExpRK2 may be preferred for moderate step sizes where its superior accuracy is most evident.

### v2

@@ 5. Examples
To conclude our manuscript, we present three examples useful for demonstrating the functionalities of order 2 ExpRK method (ExpRK2) and their comparison with other numerical methods of equal order, namely the standard order 2 RK method (RK2) and the classic order 2 Rosenbrock method (Rb2). In this way, we aim to provide both a more complete description of the numerical methods analyzed in this manuscript and a useful tool for approaching ExpRK methods. We remark that, in 2020, Montanelli and Bootland [montanelli2020solving] published a comprehensive review of exponential methods, comparing various formulas for solving stiff PDEs in multiple dimensions. Their analysis, focused on periodic boundary conditions and semilinear stiff PDEs, concluded that the ETDRK4 scheme by Cox and Matthews [cox2002exponential] (denoted as ExpRK4 in the following, for ease and uniformity of notation) remains one of the most effective approaches despite the proliferation of more complex alternatives. Their work is particularly notable for its examination of both diffusive problems (characterized by large negative eigenvalues) and dispersive problems (distinguished by their highly oscillatory nature), as well as its discussion of computational techniques for the $\varphi$-functions based on contour integrals. In line with Montanelli and Bootland’s review [montanelli2020solving], we restrict our attention to explicit exponential methods. These methods are by far the most widely employed in practice: they often yield superior performance when stiffness is predominantly carried by the linear dynamics while the nonlinear contribution remains comparatively mild. Moreover, they avoid the extra complications associated with Newton-type iterations, such as the need to solve nonlinear systems at every step and to tune iterative solver tolerances; these steps bring additional computational cost and potential convergence issues. For these reasons, explicit exponential formulations are a natural and pragmatic choice in many applications. Nevertheless, in this work we also deliberately test them on less typical, small-scale examples to probe their behavior in unconventional settings.
The three didactic examples we propose have dimension 1, 2 and 50 (where here the ‘dimension’ refers to the size of the system of ODEs), respectively. Specifically, for the 1D case, we revisit the introductory example presented by Cox and Matthews in 2002 [cox2002exponential, Sect. 4.1]; for the 2D case, we examine a classical model from nonlinear mechanics and dynamical systems, namely the Duffing oscillator [^12] For a comprehensive analysis of the nonlinear Duffing oscillator and its applications, we recommend the following volume [DuffingBook] to the interested reader.; the last example is the result of applying the pseudospectral discretization method [bdgsv16] to an Ikeda DDE [ELLG04]. In what follows, we briefly introduce three order $2$ methods and three order $4$ methods for solving stiff systems of ODEs. As already noticed, stiff ODEs require specialized numerical techniques due to their rapidly changing components existing alongside slowly evolving ones. The six methods under consideration–RK2, ExpRK2, Rb2, RK4, ExpRK4 and Rb4–represent different approaches to handling stiffness, while all maintaining second-order accuracy. Their properties suggest different performance characteristics when applied to stiff problems.
@@
In the first example we present, there exists an analytical expression for the solution $u$ of the one-dimensional ODE, and we use this exact expression in the formula above. In the second two-dimensional example, such an explicit, exact expression does not exist, and we compare the methods against a reference expression for the solution $u$. This is also the case for the third example, for which we compute the reference solution using a built-in MATLAB integrator for DDEs. Additionally, we measure the computational time required by each method, enabling a comprehensive comparison of both accuracy and efficiency. This dual assessment, as emphasized earlier, is particularly important when dealing with stiff systems, where the choice of numerical method significantly impacts the simulation’s precision and performance. Through this analysis, we aim to further illustrate the practical considerations discussed in previous sections and provide additional insights into the behavior of these methods across different problem types and stiffness regimes. All simulations were performed in MATLAB, version R2024a, on a 64-bit Microsoft Windows system.
The Six Methods
Here, we describe the six numerical methods which we have used for our simulations, namely for solving (in one case non-autonomous, in the other cases autonomous) systems of ODEs of the form:
@@
In particular, we consider three methods of order 2 and three methods of order 4.
@@
$\mathbf{a}$ $=\mathrm{e}^{hA}\mathbf{u}_{n}+h\varphi_{1}(hA)\mathbf{g}(t_{n},\mathbf{u}_{n})\,$
@@
is the Jacobian matrix, and $\gamma$ is a stabilization parameter; in particular, $\gamma=0.5$ provides a method of order 2.
Explicit Runge-Kutta 4 Method (RK4)
The classical order 4 method that we consider is RK4, which we already introduced in Section 2.2.
Exponential Runge-Kutta 4 Method (ExpRK4)
Among the order 4 exponential methods, we consider ExpRK4 [cox2002exponential, eq. (26)-(29)]. Considering again the equation,
(29) $\frac{d\mathbf{u}}{dt}=A\mathbf{u}+\mathbf{g}(t,\mathbf{u})\,$,
ExpRK4 updates the solution as follows:
$\mathbf{g}_{1}$ $=\mathbf{g}(t_{n},\mathbf{u}_{n}),\,$
$\mathbf{a}$ $=\mathrm{e}^{hA/2}\mathbf{u}_{n}+\frac{h}{2}\varphi_{1}(hA/2)\mathbf{g}_{1},\,$
$\mathbf{g}_{2}$ $=\mathbf{g(t_{n}+h/2,\mathbf{a}}),\,$
$\mathbf{b}$ $=\mathrm{e}^{hA/2}\mathbf{u}_{n}+\frac{h}{2}\varphi_{1}(hA/2)\mathbf{g}_{2},\,$
$\mathbf{g}_{3}$ $=\mathbf{g(t_{n}+h/2,\mathbf{b}}),\,$
$\mathbf{c}$ $=\mathrm{e}^{hA/2}\mathbf{a}+\frac{h}{2}(2\mathbf{g}_{3}-\mathbf{g}_{1}),\,$
$\mathbf{g}_{4}$ $=\mathbf{g}(t_{n}+h,\mathbf{c})),\,$
$b_{1}(hA)$ $=\varphi_{1}(hA)-3\varphi_{2}(hA)+4\varphi_{3}(hA),\,$
$b_{2}(hA)$ $=2\varphi_{2}(hA)-4\varphi_{3}(hA),\,$
$b_{3}(hA)$ $=-\varphi_{2}(hA)+4\varphi_{3}(hA),\,$
$\mathbf{u}_{n+1}$ $=\mathrm{e}^{hA}\mathbf{u}_{n}+h(b_{1}(hA)\mathbf{g}_{1}+b_{2}(hA)(\mathbf{g}_{2}+\mathbf{g}_{3})+b_{3}(hA)\mathbf{g}_{4}),\,$
where $\varphi_{1}$, $\varphi_{2}$ and $\varphi_{3}$ are defined from (4).
Rosenbrock 4 Method (Rb4)
The order 4 Rosenbrock-Wanner method we consider is defined by four stages. Introducing for ease of notation $f(t_{n},\mathbf{u}_{n}):=f(\mathbf{u}_{n})$, the update rule is given by:
$\left(I+\gamma_{ii}h\mathbf{J}_{n}\left(\mathbf{u}_{n}\right)\right)\mathbf{k}_{i}$
$=hf\left(\mathbf{u}_{n}+\sum_{j=1}^{i-1}\alpha_{ij}\mathbf{k}_{j}\right)+h\mathbf{J}_{n}(\mathbf{u}_{n})\sum_{j=1}^{i-1}\gamma_{ij}\mathbf{k}_{j},\quad i=1,\ldots 4,$
$\mathbf{u}_{n+1}$ $=\mathbf{u}_{n}+\sum_{i=1}^{4}\beta_{i}\mathbf{k}_{i},$
for constant $\alpha_{ij},\,\beta_{i},\,\gamma_{ij}$ which are chosen in order to satisfy the order 4 conditions (see, e.g., [hw91, Section IV.7]).
@@
Following the same approach presented by Cox and Matthews [cox2002exponential] in their analysis of ETD methods, we consider a non-autonomous ODE model with rapid decay, characteristic of stiff systems:
@@
where we have chosen k=100k=-100 to generate the rapid linear decay typical of stiff problems, in contrast with the $\mathcal{O}(1)$ time scale of the forcing term. This equation, while simple in form, captures the essential features of stiff systems: a fast decaying homogeneous solution combined with a (relatively) slow varying forcing term. For subsequent evaluation of the different numerical schemes, we utilize the exact analytical solution given by:
@@
The behavior of this differential equation can be divided into two distinct phases: an initial fast transient phase, during which the solution rapidly approaches the slow manifold, followed by a phase in which the solution evolves along this slow manifold. As noticed in [cox2002exponential], when $k<0$ and $|k|\gg 1$, the solution quickly approaches a slow manifold on which
@@
A robust numerical method must accurately capture both of these phases and work effectively when the time step $h$ is of the order of $1/|k|$. This test case allows us to evaluate how well our three order $2$ and three order $4$ methods handle the challenges posed by stiffness, particularly the balance between stability and accuracy when using relatively large time steps compared to the fastest time scale in the system. We examined the performance of the methods over the interval $0,\pi/2]$ using various time steps ranging from $h=0.0001$ to $h=0.1$, comparing both the accuracy of the numerical solutions against the analytical solution and the computational efficiency of each method. To estimate the elapsed time, we performed each simulation several times (1000 in the first test, 10 in the second and 100 in the third) and reported the average.
Table 2. Numerical methods performance for various step sizes (order 2 methods).
@@
$1\times 10^{-4}$ $1.2551\times 10^{-9}$ $8.3276\times 10^{-10}$ $2.3212\times 10^{-9}$ 0.05355 0.05829 0.05625
$5\times 10^{-4}$ $3.1432\times 10^{-8}$ $2.0438\times 10^{-8}$ $3.1446\times 10^{-8}$ 0.01183 0.01188 0.01070
$1\times 10^{-3}$ $1.2779\times 10^{-7}$ $8.0986\times 10^{-8}$ $1.6274\times 10^{-7}$ 0.00556 0.00584 0.00585
$5\times 10^{-3}$ $3.8530\times 10^{-6}$ $1.9199\times 10^{-6}$ $4.4691\times 10^{-6}$ 0.00118 0.00131 0.00118
$1\times 10^{-2}$ $2.3082\times 10^{-5}$ $7.5741\times 10^{-6}$ $2.1139\times 10^{-5}$ 0.00059 0.00065 0.00061
$5\times 10^{-2}$ $7.0246\times 10^{30}$ $4.8629\times 10^{-5}$ $1.4075\times 10^{-4}$ 0.00010 0.00013 0.00011
$1\times 10^{-1}$ $2.9523\times 10^{27}$ $2.5459\times 10^{-4}$ $4.1347\times 10^{-2}$ 0.00005 0.00006 0.00005
Table 3. Numerical methods performance for various step sizes (order 4 methods).
$h$ RK4 Error ExpRK4 Error Rb4 Error RK4 Time (s) ExpRK4 Time (s) Rb4 Time (s)
$1\times 10^{-4}$ $1.0583\times 10^{-14}$ $4.4262\times 10^{-10}$ $7.3560\times 10^{-14}$ 0.05772 0.07283 0.17238
$5\times 10^{-4}$ $6.5026\times 10^{-12}$ $1.9799\times 10^{-12}$ $4.2579\times 10^{-11}$ 0.01216 0.01484 0.03288
$1\times 10^{-3}$ $1.0440\times 10^{-10}$ $8.0308\times 10^{-13}$ $6.2264\times 10^{-10}$ 0.00576 0.00692 0.01561
$5\times 10^{-3}$ $7.6433\times 10^{-8}$ $6.7939\times 10^{-13}$ $8.0284\times 10^{-6}$ 0.00130 0.00150 0.00327
$1\times 10^{-2}$ $5387\times 10^{-6} 1.1$$1421\times 10^{-11}$ $5.7421\times 10^{-6}$ 0.00061 0.00073 0.00161
$5\times 10^{-2}$ $6.4073\times 10^{36}$ $3.0104\times 10^{-9}$ $1.7416\times 10^{-4}$ 0.00011 0.00013 0.00029
$1\times 10^{-1}$ $5.8631\times 10^{40}$ $4.7821\times 10^{-8}$ $5.2099\times 10^{-4}$ 0.00006 0.00007 0.00016
@@
Figure 1. Relative error as a function of step size, showing how accuracy of each method of order 2 (1(a)) and order 4 (1(c)) evolves with different step choices and relative error as a function of computational time, highlighting the efficiency of methods of order 2 (1(b)) and order 4 (1(d))
@@
The results presented in Figure 1 and Tables 2-3 highlight the distinct behaviors of the three order $2$ and the three order $4$ (respectively) numerical methods under analysis.
@@
RK2 and RK4 demonstrate stability only for small step sizes, specifically for values of $h\leq 0.01$. However, for $h\geq 0.05$, the relative error grows uncontrollably. This behavior confirms a strict stability limit around $h=0.01$, beyond which the method fails to provide reliable results. On the other hand, both exponential methods and Rb4 exhibit excellent stability across all tested step sizes. Even for larger ones, the methods maintain a reasonably small relative error. Rb2 also demonstrates stable behavior, albeit within a more restricted range. For step sizes up to $0.01$, the method produces accurate results. However, for $h=0.05$, the relative error starts increasing significantly, and for $h=0.1$ it becomes quite large ($1.53\times 10^{-1}$). The superior accuracy makes ExpRK2 a strong candidate for solving stiff problems among those of order 2. With respect to those of order 4, ExpRK4 provides much more accurate results than its Rosenbrock counterpart for most values of $h$, although Rb4 behaves well for all tested step sizes.
Regarding computational efficiency, differences in computation times among the methods of order 2 are negligible, making observations about performance insignificant. In contrast, ExpRK4 appears to be the most efficient among the methods of order 4. It is worth noting that a relative error close to machine precision is already reached using $h=0.005$, which makes the use of smaller step sizes not sensible for this example.
@@
Considering both accuracy and efficiency, ExpRK2 emerges as the best-balanced method among those of order 2. It sustains high accuracy even with large step sizes, allowing for significantly larger steps compared to the other methods. Although it has a slight computational overhead per step, the reduced number of required steps makes it more efficient overall.
Rb2 offers similar computation times for very small step sizes but with stricter stability constraints: the method appears to work very effectively for a problem of small scale. RK2, in contrast, proves to be the least suitable method for this stiff problem, as it requires extremely small time steps to maintain stability and lacks competitiveness in both accuracy and computational efficiency.
These findings align with the conclusions drawn by Cox and Matthews in [cox2002exponential], confirming the superiority of exponential methods for solving stiff problems. In particular, ExpRK2 offers a significant advantage in terms of stability and accuracy, permitting the use of much larger step sizes than conventional methods. The superiority of ExpRK4 with respect to the tested methods of the same order is even more evident. For stiff problems such as the one studied here, characterized by a parameter k=100k=-100, exponential methods represent the optimal choice among the three methods of the corresponding order, striking an ideal balance between accuracy, stability, and computational efficiency.
@@
Table 4. Numerical methods performance for various step sizes (order 2 methods).
$1\times 10^{-4}$ $6.2072\times 10^{-5}$ $6.0105\times 10^{-6}$ $8.7547\times 10^{-5}$ 0.09026 0.54986 1.13880
$5\times 10^{-4}$ $1.9363\times 10^{-3}$ $5767\times 10^{-4} 2.2$$2968\times 10^{-3}$ 0.01701 0.09588 0.20080
$1\times 10^{-3}$ $9.6628\times 10^{-3}$ $1.5697\times 10^{-3}$ $9.7327\times 10^{-3}$ 0.00866 0.04709 0.10134
$5\times 10^{-3}$ $6.3873\times 10^{-1}$ $1.4548\times 10^{-1}$ $3.4409\times 10^{-1}$ 0.00176 0.00874 0.01715
$1\times 10^{-2}$ $2.9671\times 10^{0}$ $9.5176\times 10^{-1}$ $1.1835\times 10^{0}$ 0.00092 0.00674 0.01204
$5\times 10^{-2}$ NaN $1.3049\times 10^{0}$ $1.2752\times 10^{-1}$ 0.00008 0.00224 0.00277
$1\times 10^{-1}$ NaN $1.6391\times 10^{0}$ $8.3403\times 10^{-1}$ 0.00002 0.00059 0.00113
Table 5. Numerical methods performance for various step sizes (order 4 methods).
$h$ RK4 Error ExpRK4 Error Rb4 Error RK4 Time (s) ExpRK4 Time (s) Rb4 Time (s)
$1\times 10^{-4}$ $4.7521\times 10^{-12}$ $1.0584\times 10^{-9}$ $6.0232\times 10^{-11}$ 0.83861 15.54832 2.15542
$5\times 10^{-4}$ $6.7006\times 10^{-9}$ $3246\times 10^{-9} 2.2$$2207\times 10^{-8}$ 0.16769 0.80160 0.50081
$1\times 10^{-3}$ $1.3075\times 10^{-7}$ $3.8983\times 10^{-8}$ $2.0253\times 10^{-7}$ 0.06966 0.20216 0.20446
$5\times 10^{-3}$ $1.9809\times 10^{-4}$ $4.9790\times 10^{-5}$ $6.3571\times 10^{-4}$ 0.01447 0.02514 0.04033
$1\times 10^{-2}$ $5.4686\times 10^{-3}$ $1.3051\times 10^{-3}$ $2.5171\times 10^{-2}$ 0.00977 0.01663 0.02508
$5\times 10^{-2}$ $4.1232\times 10^{-1}$ $3.2182\times 10^{0}$ $4.6251\times 10^{-1}$ 0.00217 0.00512 0.00734
$1\times 10^{-1}$ $1.9771\times 10^{0}$ NaN $3.2802\times 10^{-2}$ 0.00105 0.00300 0.00378
Figure 2. Relative error as a function of step size, showing how accuracy of each method of order 2 (2(a)) and order 4 (2(c)) evolves with different step choices and relative error as a function of computational time, highlighting the efficiency of methods of order 2 (2(b)) and order 4 (2(d))
The results presented in Figure 2 and Tables 4-5 provide a comprehensive analysis of six numerical methods of order 2 and 4 applied to the Duffing oscillator system [^13] All relative errors are computed with respect to a high-precision numerical solution obtained using MATLAB’s ode15s solver, as no analytical solution is available for this nonlinear system. with $\omega=1$ and $k=100$.
RK2 demonstrates severe stability limitations when applied to this problem. For step sizes $h\leq 0.001$, the method produces reasonable errors, but as $h$ increases to $0.01$, the relative error grows to approximately $2.97$ and, most importantly, for larger step sizes ($h\geq 0.05$), the method fails entirely, producing NaN values. On the other hand, stability constraints on RK4 are not as evident. This suggests that the problem, with these parameters, can be considered only midly stiff.
@@
ExpRK2 demonstrates superior accuracy for most step sizes up to $h=0.01$, particularly excelling at moderate step sizes where it maintains errors significantly lower than the other methods. For $h=5\times 10^{-4}$ and $h=1\times 10^{-3}$, ExpRK2 shows errors approximately an order of magnitude smaller than both RK2 and Rb2, despite not being designed so as to capture stiffness phenomena caused by the nonlinear terms. However, for larger step sizes ($h\geq 0.05$), the errors increase substantially, exceeding the error of Rb2 at these larger step sizes. ExpRK4 exhibits similar behavior, in that it provides more accurate results for the intermediate step sizes. However, the difference in this case is not as significant.
Regarding computational efficiency, across all step sizes small enough for RK2 to obtain a solution, RK2 is the fastest method. ExpRK2 occupies an intermediate position, with computation times around 3 times higher than RK2 for a given (relative) error. Rb2 emerges as the most computationally intensive method, requiring instead around 10 times the time required by RK2 for a given error. Among the methods of order 4, the classical RK is the fastest again, but the exponential method ExpRK4 is faster than the Rosenbrock method of the same order for intermediate step sizes. In any case, the difference among the three methods in terms of performance is not as significant as within the methods of order 2.
@@
The analysis of these numerical methods for the Duffing oscillator reveals important trade-offs between accuracy, stability, and computational efficiency. RK2, while computationally inexpensive, suffers from severe stability limitations that render it unsuitable for step sizes larger than $0.001$. Rb2 is the most accurate for step sizes larger than $0.01$, while it becomes less accurate than ExpRK2 for smaller step sizes. Its non-monotonic error behavior at larger step sizes warrants further investigation but could potentially be exploited in specific applications. As for the methods of order 4, the Rosenbrock one appears again to be the most accurate for larger step sizes, while ExpRK4 may be a better choice for applications with higher accuracy constraints.
ExpRK2 demonstrates excellent accuracy, particularly at larger step sizes where it significantly outperforms both RK2 and Rb2. Indeed, it might be the best combination in terms of accuracy and efficiency, depending on the specific application. The same can be inferred about ExpRK, when compared to the other methods of order 4.
For the Duffing oscillator with the given parameters, the exponential methods appear to offer the most favorable combination of stability, accuracy, and computational efficiency among the methods of the same order. They maintain acceptable accuracy even at large step sizes and does so with reasonable computational cost. This makes them particularly suitable for applications where a balance between accuracy and efficiency is required. However, in applications where low latency is of outmost importance, the Rosenbrock methods may be preferred for moderate step sizes where its superior accuracy is most evident.